# An Entropy-based Text Watermarking Detection Method

### v1

@@ 6.1 Experiment Settings
Tasks and Datasets. First of all, we used the KGW Kirchenbauer et al. (2023a) for watermark generation in all the following tasks.
@@
Besides code generation, we fine-tuned a T5-small Raffel et al. (2023) model for data-to-text generation. Specifically, we utilized Wu et al. (2022)’s version of the Rotowire dataset Wiseman et al. (2017), which was converted for a text-to-table task. The dataset consists of NBA game records, presented as tables, and human-written descriptions. We used the table information as input and asked the fine-tuned T5 to generate descriptive texts by binomial sampling and beam-search. For each input prompt, the models are expected to generate 200 ± 5 tokens.
@@
Baselines and Evaluation Metrics. We selected two detection methods as the baselines to compare with. The first is KGW Kirchenbauer et al. (2023a), which is the original detector. The second is SWEET Lee et al. (2023), which excludes low-entropy tokens during detection. We follow the instructions in SWEET to select entropy threshold, namely from the median, mean, and 3rd quartile (Q3) values of the entropy computed from human texts. For evaluation, we set the false positive rate (FPR) under 5%, following Zhao et al. (2023)’s instruction. We reported the true positive rate (TPR) and F1 score for each method.
Hyper-parameters. For watermarked generation, we set $\gamma$ and $\delta$ to 0.5 and 2 respectively. In text generation using binomial sampling, we set the sampling temperature to 0.7. In text generation using beam-search, we set the number of beams to 8. In order to avoid excessive repetition of text, we set "no_repeated_ngrams" to 16.
@@ 6.2 Main Results
Before detection, we measured the generation quality of all generated texts. Specifically, we are able to achieve 31.0% and 30.4% in pass@1Chen et al. (2021) for HumanEval and MBPP dataset respectively, comparable to Lee et al. (2023)’s. For the Rotowire dataset, our generated output can achieve 7.24 in text perplexity (PPL). For the C4 dataset, our generated output can achieve 3.25 in PPL. PPL is calculated using OPT-6.7B Zhang et al. (2022).
@@ 6.3 Analysis
Impact of Entropy-based Weights. In figure 2, we give a few visualizations to show the reason why our method surpassed other baselines in detection performance. Figure 2(a) shows the z-scores of both watermarked and human texts in different datasets, detected by different methods. Our detection methods would result in overall higher z-score for watermarked texts, and slightly lower z-score for human texts. The z-score increase is more significant in low-entropy code generation tasks than in the Rotowire dataset. By enlarging the gap between the z-scores of watermarked and human texts, our detector could better distinguish the watermarked texts. In figure 2(b), we plotted the relationship of entropy-based weight and probability of being watermarked. In watermarked text, the probability is proportional to the token weight. That is to say, the tokens assigned with a large weight are very likely to be watermarked and contribute to a larger z-score. On the contrary, in human-written texts, there doesn’t exist such a relationship, explaining why the z-scores of human text remain low.
@@
Compared to the KGW detection, both SWEET and our proposed EWD would demand extra time to compute token entropy. Specifically, our detector would also compute weight given the token entropy before calculating the final z-score. According to table 2, the detection time per text is 0.0391, 0.0830 and 0.0831 seconds for KGW, SWEET and our detector respectively. In spite of being double the value of KGW, our detector remains highly efficient, and the difference with SWEET is virtually insignificant.

### v2

@@ 6.1 Experiment Settings
Tasks and Datasets. First of all, we used the KGW Kirchenbauer et al. (2023a) to embed watermark in all the following tasks.
@@
Besides, we fine-tuned a T5-small Raffel et al. (2023) model for data-to-text generation. Specifically, we utilized Wu et al. (2022)’s version of the Rotowire dataset Wiseman et al. (2017), which was converted for a text-to-table task. The dataset consists of NBA game records, presented as tables, and human-written descriptions. We used the table information as input and generate descriptive texts. For each input prompt, the models are expected to generate 200 ± 5 tokens.
@@
Baselines and Evaluation Metrics. We selected two detection methods as baselines. The first is KGW detector Kirchenbauer et al. (2023a). The second is SWEET’s detector Lee et al. (2023), which excludes low-entropy tokens during detection. We follow the instructions in SWEET to select entropy threshold from the median, mean, and 3rd quartile (Q3) values of the entropy computed from human texts. For evaluation, we set the false positive rate (FPR) under 5%, following Zhao et al. (2023)’s instruction. We reported the true positive rate (TPR) and F1 score for each method.
Hyper-parameters. For watermarked algorithm, we set $\gamma$ and $\delta$ to 0.5 and 2 respectively. Binomial sampling set the sampling temperature to 0.7. Beam-search set the number of beams to 8. To avoid excessive repetition of text, we set "no_repeated_ngrams" to 16.
@@ 6.2 Main Results
Before detection, we measured the generation quality of all generated texts. Specifically, we are able to achieve 31.0% and 30.4% in pass@1Chen et al. (2021) for HumanEval and MBPP dataset respectively, comparable to Lee et al. (2023)’s. For the Rotowire and C4 dataset, our generated output can achieve 7.24 and 3.25 in text perplexity (PPL) using OPT-6.7B Zhang et al. (2022).
@@ 6.3 Analysis
Impact of Entropy-based Weights. In figure 2, we gave a few visualizations to show the reason why EWD surpassed other baselines in detection performance. Figure 2(a) shows the z-scores of both watermarked and human texts in different datasets, detected by different methods. Our EWD method would result in overall higher z-score for watermarked texts, and slightly lower z-score for human texts. The z-score increase is more significant in low-entropy code generation tasks than in the Rotowire dataset. By enlarging the gap between the z-scores of watermarked and human texts, our detector could better distinguish the watermarked texts. In figure 2(b), we plotted the relationship of a token’s entropy-based weight and probability of that token being watermarked. In watermarked text, the probability is proportional to the token weight. That is to say, the tokens assigned with a large weight are very likely to be watermarked and contribute to a larger z-score. On the contrary, in human-written texts, there doesn’t exist such a relationship, explaining why the z-scores of human text remain low.
@@
Compared to the KGW detection, both SWEET and our proposed EWD would demand extra time to compute token entropy. Specifically, our detector would also compute weight given the token entropy before calculating the final z-score. According to table 2, the detection time per text is 0.0391, 0.0830 and 0.0831 seconds for KGW, SWEET and our EWD respectively. In spite of being double the value of KGW, our detector remains highly efficient, and the difference with SWEET is virtually insignificant.
Performance against Attacks.
Table 3: Detection performance of back-translated watermarked texts using different detection methods.
C4 w/ Back-translation
Methods 1%FPR 5%FPR
TPR F1 TPR F1
KGW 0.894 0.940 0.950 0.952
SWEET 0.890 0.938 0.930 0.940
EWD 0.900 0.943 0.942 0.947
Watermarked texts usually would be edited, which would cause detection performance drop. Here we also evaluated the baseline methods and our EWD under this setting by utilizing back-translation as the attack to remove watermark. Specifically, we firstly generated watermarked texts using the C4 dataset, then translated them from English to French, and later back to English again for detection. In table 3, we can observe that our EWD would achieve the best detection accuracy under 1% FPR and similar results with KGW under 5% FPR.